OpenStudy (anonymous):

simplify (x - 2)(x + 2) = 5^1

OpenStudy (hba):

Is it, \[\huge\ (x-2)(x+2)=5^1\]

OpenStudy (hba):

@ma12209 ?

OpenStudy (anonymous):

x^2+2x-2x-4=5^1 x^2-4=5 x^2=5+4 x^2=9 x=3 i think hehe

OpenStudy (hba):

@miszery Giving direct answers is not allowed on OS. Please let the asker participate. If you are still not sure please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

thanks :) everyone

OpenStudy (hba):

@miszery Is that understood ?

OpenStudy (hba):

@skullpatrol Not you dude :P

OpenStudy (hba):

@ma12209 You should check this \[\sqrt{9}=\pm3\]

OpenStudy (anonymous):

sorry I don't understand I didn;t know the answer I was trying to figure it out

OpenStudy (hba):

@miszery I hope you don't do the same think the next time.

OpenStudy (anonymous):

sorry you guys for the problem i caused.....thanks for your help though ...i think its better if i just dont ask

OpenStudy (hba):

@ma12209 Well the point is that you should participate in the solution :) Nothing else.

OpenStudy (skullpatrol):

"Ask" but don't demand the answer...

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok sorry and thanks

OpenStudy (hba):

Welcome :)

OpenStudy (skullpatrol):

Remember to try and get involved in the solution process.

OpenStudy (jiteshmeghwal9):

(x-2)(x+2)=5^1 Use the identity \((a+b)(a-b)=a^2-b^2\) x^2-4=5 x^2=5+4=9 take square root both sides \(\sqrt{x^2}=\sqrt{9}\)

OpenStudy (jiteshmeghwal9):

now solve this for x :)